### v2

@@ 4.2.3 Detection-only Datasets (Camera-only)
RoScenes [zhu2024roscenes] and WIBAM [howe2021weakly] datasets rely on monocular or BEV camera views, designed for object detection and scene understanding without additional 3D sensing or tracking. The RoScenes dataset, released in 2024, aims to enhance vision-based Bird’s Eye View (BEV) approaches using data from 14 highway scenes in China [zhu2024roscenes]. It includes 1.21 million images and 21.3 million 3D bounding box labels covering 64,000m². The sensor setup features 6–12 roadside cameras mounted on 10m poles, two cameras (different zooms) on each side, and drones flying at 300m to capture full highway coverage without blind spots. Its strengths include a massive scale of data, diverse viewpoints that reduce occlusions, and an efficient BEV-to-3D annotation method. However, the dataset is limited to highway scenarios under sunny conditions, lacking scenes like tunnels or intersections. The WIBAM (Wide Baseline Multiview) dataset [howe2021weakly], released in October 2021, supports weakly supervised training of monocular 3D object detectors using traffic surveillance footage from elevated cameras at UK intersections. It addresses challenges in estimating 3D vehicle poses from elevated views. The dataset includes 33,000 images captured by four unsynchronized RGB cameras (originally 2560×1140, downsampled to 1920×1080), with 116,702 automatically generated 2D training annotations and 1,651 manually labeled 3D bounding boxes in the test set. It enables pose prediction and employs a weak supervision approach using multi-view reprojection loss. Strengths include reduced labeling costs and real-world applicability; however, its fixed elevated viewpoint limits use in autonomous driving, and auto-generated annotations may introduce noise.
@@ 4.3 Vehicle-to-Language (V2L) Datasets
[ishaq2025drivelmm] DriveLMM.o1 2025 NuScenes Dataset +1.9K 18K Yes Yes Git
[chen2025automated] CODA-LM 2025 CODA Dataset +9K 63K Yes No Git
[chiu2025v2v] V2V-QA 2025 V2V4Real, V2X-Real +48K 1.45M Yes Yes Git
[sima2024drivelm] DriveLM 2024 NuScenes, Carla +74K 375K Yes Yes Git
[inoue2024nuscenes] NuScenes-MQA 2024 NuScenes Dataset +34K 1.5M Yes No Git
[wang2025omnidrive] OmniDrive 2024 NuScenes, OpenLane-v2 +34K 450K Yes No Git
[tian2024tokenize] TOKEN 2024 NuScenes, DriveLM +28K 434K Yes No N/A
[cao2024maplm] MAPLM-QA 2024 Original Dataset +14K 61K Yes Yes Git
[marcu2024lingoqa] Lingo-QA 2024 Original Dataset +28K 420K Yes No Git
[qian2024nuscenes] NuScenes-QA 2024 NuScenes Dataset +34K 460K Yes Yes Git
[ding2024holistic] NuInstruct 2024 NuScenes Dataset +11K 91K Yes No Git
[malla2023drama] DARAMA 2023 Original Dataset +18K 103K Yes No Web
@@ 4.3.1 Camera + LiDAR Datasets
DriveLMM.o1 [ishaq2025drivelmm], V2V-QA [chiu2025v2v], MAPLM-QA [cao2024maplm], NuScenes-QA [qian2024nuscenes], and DriveLM [sima2024drivelm] datasets provide both RGB imagery and point cloud data, making them suitable for spatially-grounded language tasks, 3D-aware VQA, and multi-sensor fusion in perception and reasoning. The DriveLMM.o1 dataset [ishaq2025drivelmm], released in March 2025 by Mohamed Bin Zayed University of AI, the DriveLMM-o1 dataset serves as a multimodal benchmark to evaluate LLM reasoning in autonomous driving. It includes over 18,000 visual question answering (VQA) samples across four subsets, covering diverse real-world driving scenarios that involve perception, prediction, and planning tasks. Each sample contains synchronized multiview images, LiDAR data, temporal cues, questions, ground-truth answers, and structured step-by-step reasoning. Designed to test LLMs’ interpretability and multi-step decision-making, the dataset also provides tools for data parsing and custom evaluation metrics. It supports research in visual reasoning, sensor fusion, and end-to-end autonomous decision-making. Strengths include detailed annotations, dynamic scene diversity, and structured explanations. Limitations include limited geographic and environmental diversity, as well as high computational demands due to data complexity. DriveLMM-o1 sets a new standard for evaluating multimodal LLM reasoning in Avs.
@@
The V2V-QA dataset [chiu2025v2v], released in April 2025 by Carnegie Mellon University and NVIDIA, is a large-scale multimodal benchmark for enhancing collaborative autonomous driving via language-based reasoning. Built on real-world V2V4Real and V2X-Real datasets, it includes over 48,000 frames and 1.45 million question-answer (QA) pairs that simulate vehicle-to-vehicle communication. The QA system supports object grounding, identification, and driving planning tasks in multi-agent environments using synchronized LiDAR and camera data. Each frame contains five QA types generated through geometric templates and contextual rules. The dataset enables training and evaluation of vision-language models, like V2V-LLM, which outperform standard LLMs in cooperative perception and intention prediction. Strengths include its focus on explainability, situational awareness, and multi-agent interaction. Limitations include dependency on specific sensor setups, restricted geographic coverage, high computational demands, and uncertain usage permissions. NuScenes-QA dataset released in 2023 by Fudan University [qian2024nuscenes]. It is a large-scale VQA benchmark built on the nuScenes dataset to enhance reasoning in autonomous driving. It contains 460,000 QA pairs derived from 34,000 driving scenes with synchronized RGB images and LiDAR point clouds. Using 3D scene graphs and logic-based templates adapted from the CLEVR benchmark, researchers generated diverse questions ranging from basic attributes to complex spatiotemporal inference. Temporal questions require multi-frame reasoning to track changes and object motion. The dataset evaluates vision-language models and includes baseline results from Transformer-based models. Although NuScenes-QA advances VQA for driving tasks, limitations include a lack of background scene understanding, limited linguistic diversity due to templated questions, and shallow semantic depth.
@@
The MAPLM-QA dataset [cao2024maplm], introduced in 2024 through a collaboration between Tencent’s T Lab and several universities. It is a large-scale multimodal benchmark for vision-language reasoning in autonomous driving. It includes 61,000 QA pairs linked to 14,000 sensor-rich frames featuring panoramic RGB images, 3D LiDAR point clouds, and HD maps from diverse real-world environments like intersections, highways, and rural roads. To overcome limitations of prior VQA datasets, MAPLM-QA challenges LLMs to reason about road layouts, markings, and navigation cues. Data was collected using vehicles equipped with six cameras, LiDAR, and GPS/IMU, with QA annotations refined through a vision model and human verification. The dataset connects language prompts to spatial and environmental features and offers JSON-formatted tools for integration and evaluation. By evaluating LLMs like CLIP and LLaMA, this benchmark highlights their effectiveness in real-world grounding and exposes their limitations in spatial reasoning. Although it is limited by geographic scope and high computational requirements, MAPLM-QA is a pioneering resource for advancing LLM-based perception and planning. The DriveLM dataset [sima2024drivelm], released in January 2025 by OpenDriveLab, is a large-scale benchmark for evaluating LLMs and VLMs in autonomous driving via Graph VQA. Unlike standard perception-control datasets, DriveLM emphasizes multi-step reasoning across perception, prediction, planning, and behavior tasks. It includes over 74,000 annotated frames across two subsets: DriveLM-nuScenes (4,871 frames, $\sim$91.4 QAs/frame) and DriveLM-CARLA (70,006 frames, $\sim$20.5 QAs/frame). DriveLM-nuScenes uses keyframe-based manual annotation with automatic QA generation, while DriveLM-CARLA relies on expert-driven simulation and sensor data (RGB + LiDAR). DriveLM’s strengths lie in explainability, zero-shot generalization, and structured decision flow. Limitations include reliance on simulation-based data and the high resource demands.
@@ 4.3.2 Camera-only Datasets
CODA-LM [chen2025automated], NuInstruct [ding2024holistic], OmniDrive [wang2025omnidrive], Lingo-QA [marcu2024lingoqa], and DARAMA [malla2023drama] datasets rely solely on RGB visual input, often from front-facing or panoramic cameras, and are commonly used in visual-language reasoning, question answering, and instruction-following without explicit 3D spatial data. Dalian University of Technology introduced the CODA-LM dataset [chen2025automated] in 2024 as a comprehensive benchmark for assessing Large Vision-Language Models (LVLMs) in practical autonomous driving systems. In order to evaluate multimodal understanding, it consists of 9,768 high-resolution scenes with uncommon driving events combined with textual prompts. CODA-LM supports three subtasks: general perception (object/event detection), regional perception (localized object description), and driving suggestion (context-based decision reasoning). It uses a hierarchical evaluation system with automated text-only LLM scoring, which offers better alignment with human judgment than multimodal evaluators. Additionally, it introduces CODA-VLM, a fine-tuned model that outperforms all open-sourced driving LVLMs and achieves GPT-4V-level results. Strengths include its focus on safety-critical corner cases, structured reasoning, and reduced reliance on human evaluation. Limitations include restricted coverage of all rare driving scenarios and high computational demands, which may limit accessibility.
@@
In 2024, HKUST, Sun Yat-Sen University, and Huawei Noah’s Ark Lab collaborated to release the NuInstruct dataset [ding2024holistic]. It is a large-scale instruction-based multimodal benchmark for autonomous driving. It contains more than 91,000 instruction-response pairs derived from multiview driving videos, addressing limitations in prior datasets by incorporating spatial-temporal data and diverse perspectives. The dataset is organized into 17 subtasks across four categories, including perception, prediction, risk, and planning, with reasoning to support real-world driving decision-making processes. QA pairs are generated using an SQL-based method and require reasoning across frames and views. Strengths include comprehensive coverage of driving tasks and temporal Bird’s-Eye-View cross-view reasoning. However, it lacks 3D object detection and traffic light tasks, requires high computational resources, and may need domain adaptation for generalization. OmniDrive dataset [wang2025omnidrive], released in April 2025 by Hong Kong Polytechnic University and Beijing Institute of Technology. It is a state-of-the-art multimodal dataset that links 3D driving tasks with LLMs using counterfactual reasoning. Built on nuScenes, it adds 34,000 annotated frames and 434,000 QAs focused on ”what if” scenarios that bridge perception, planning, and reasoning. The dataset includes factual and counterfactual QA pairs, trajectory data, and keyframes in JSON format for transformer-based pipelines. Two evaluation agents, Omni-Q (trajectory supervision) and Omni-L (vision-language alignment), show improved scene understanding and planning response. OmniDrive’s strengths lie in integrating perception and decision-making while enhancing interpretability and explainability for AVs. Limitations include synthetic scene generation that reduce realism and high computational demands.
@@
The Lingo-QA dataset [marcu2024lingoqa], released in September 2024 by Wayve Technologies, is a large-scale multimodal benchmark for evaluating LLMs in autonomous driving. It includes 28,000 short video scenarios paired with 419,000 QA pairs covering tasks like hazard anticipation, traffic comprehension, and high-level route reasoning. Lingo-QA is built for fine-grained evaluation with QA annotations to evaluate visual perception and decision-making. Key strengths of the Lingo-QA dataset include robust annotation, multi-level reasoning, and standardized assessment. However, its limitations include reliance on short, front-facing video clips with no LiDAR or sensor data, limited scalability (up to 7B LLMs), and Lingo-Judge’s dependence on the dataset and lack of support for diverse response formats or preference-based evaluation. DRAMA dataset (Driving Risk Assessment Mechanism with A captioning module) [malla2023drama], released in 2022 by Honda Research Institute USA, is a multimodal benchmark for developing risk-aware and explainable decision-making in autonomous vehicles. The DRAMA dataset combines video and object-level QA with free-form captions explaining driver reactions, linking risk localization to natural language reasoning. It features synchronized video from two front-facing cameras, along with CAN and IMU data. Annotations include open- and closed-ended questions about risk identification, causes, and reasoning, supported by natural language explanations. DRAMA’s key strength lies in integrating multimodal risk perception with interpretability for LVLMs. However, its reliance on subjective human annotations, limited coverage of failure cases, and the risks of deploying incorrect models in safety-critical settings require careful use and additional validation for real-world applications.
@@ 4.4 Vehicle-to-Vehicle (V2V) Datasets
[li2024multiagent] Open Mars 2024 USA +2M RGB, LiDAR, Trajectories Yes Yes No Git
[lu2024extensible] OPV2V-H 2024 Carla +9K RGB, LiDAR, Depth Yes Yes No Git
[hu2023collaboration] OPV2V+ 2023 Carla 11,464 RGB, LiDAR, Depth Yes Yes No Web
@@
[xu2023v2v4real] V2V4Real 2023 USA 60K RGB, LiDAR, HDMaps Yes Yes No Web
[axmann2023lucoop] LUCOOP 2023 Germany +54K LiDAR, IMU, GNSS Yes Yes No Web
[hu2022where2comm] CP-UAV 2022 AirSim 131.9K RGB, LiDAR, BEW Yes Yes No Git
[xu2022opv2v] OPV2V 2022 Carla 11,464 RGB, LiDAR, BEW Yes Yes No Web
[arnold2021fast] CODD 2021 Carla 8,783 LiDAR, Transform.txt No Yes No Git
[yuan2021comap] COMAP 2021 Carla +4.3K RGB, LiDAR, Seg-Mask Yes Yes No Git
[chen2019cooper] T&J Cooper 2019 USA 100 RGB, LiDAR, GPS Yes Yes Yes N/A

### v3

@@ 4.2.3 Detection-only Datasets (Camera-only)
RoScenes [323] and WIBAM [87] datasets rely on monocular or BEV camera views, designed for object detection and scene understanding without additional 3D sensing or tracking. The RoScenes dataset, released in 2024, aims to enhance vision-based Bird’s Eye View (BEV) approaches using data from 14 highway scenes in China [323]. It includes 1.21 million images and 21.3 million 3D bounding box labels covering 64,000m². The sensor setup features 6–12 roadside cameras mounted on 10m poles, two cameras (different zooms) on each side, and drones flying at 300m to capture full highway coverage without blind spots. Its strengths include a massive scale of data, diverse viewpoints that reduce occlusions, and an efficient BEV-to-3D annotation method. However, the dataset is limited to highway scenarios under sunny conditions, lacking scenes like tunnels or intersections. The WIBAM (Wide Baseline Multiview) dataset [87], released in October 2021, supports weakly supervised training of monocular 3D object detectors using traffic surveillance footage from elevated cameras at UK intersections. It addresses challenges in estimating 3D vehicle poses from elevated views. The dataset includes 33,000 images captured by four unsynchronized RGB cameras (originally 2560×1140, downsampled to 1920×1080), with 116,702 automatically generated 2D training annotations and 1,651 manually labeled 3D bounding boxes in the test set. It enables pose prediction and employs a weak supervision approach using multi-view reprojection loss. Strengths include reduced labeling costs and real-world applicability; however, its fixed elevated viewpoint limits use in autonomous driving, and auto-generated annotations may introduce noise.
@@ 4.3 Vehicle-to-Language (V2L) Datasets
[103] DriveLMM.o1 2025 NuScenes Dataset +1.9K 18K Yes Yes Git
[31] CODA-LM 2025 CODA Dataset +9K 63K Yes No Git
[50] V2V-QA 2025 V2V4Real, V2X-Real +48K 1.45M Yes Yes Git
[208] DriveLM 2024 NuScenes, Carla +74K 375K Yes Yes Git
[102] NuScenes-MQA 2024 NuScenes Dataset +34K 1.5M Yes No Git
[238] OmniDrive 2024 NuScenes, OpenLane-v2 +34K 450K Yes No Git
[218] TOKEN 2024 NuScenes, DriveLM +28K 434K Yes No N/A
[27] MAPLM-QA 2024 Original Dataset +14K 61K Yes Yes Git
[167] Lingo-QA 2024 Original Dataset +28K 420K Yes No Git
[185] NuScenes-QA 2024 NuScenes Dataset +34K 460K Yes Yes Git
[61] NuInstruct 2024 NuScenes Dataset +11K 91K Yes No Git
[163] DARAMA 2023 Original Dataset +18K 103K Yes No Web
@@ 4.3.1 Camera + LiDAR Datasets
DriveLMM.o1 [103], V2V-QA [50], MAPLM-QA [27], NuScenes-QA [185], and DriveLM [208] datasets provide both RGB imagery and point cloud data, making them suitable for spatially-grounded language tasks, 3D-aware VQA, and multi-sensor fusion in perception and reasoning. The DriveLMM.o1 dataset [103], released in March 2025 by Mohamed Bin Zayed University of AI, the DriveLMM-o1 dataset serves as a multimodal benchmark to evaluate LLM reasoning in autonomous driving. It includes over 18,000 visual question answering (VQA) samples across four subsets, covering diverse real-world driving scenarios that involve perception, prediction, and planning tasks. Each sample contains synchronized multiview images, LiDAR data, temporal cues, questions, ground-truth answers, and structured step-by-step reasoning. Designed to test LLMs’ interpretability and multi-step decision-making, the dataset also provides tools for data parsing and custom evaluation metrics. It supports research in visual reasoning, sensor fusion, and end-to-end autonomous decision-making. Strengths include detailed annotations, dynamic scene diversity, and structured explanations. Limitations include limited geographic and environmental diversity, as well as high computational demands due to data complexity. DriveLMM-o1 sets a new standard for evaluating multimodal LLM reasoning in Avs.
@@
The V2V-QA dataset [50], released in April 2025 by Carnegie Mellon University and NVIDIA, is a large-scale multimodal benchmark for enhancing collaborative autonomous driving via language-based reasoning. Built on real-world V2V4Real and V2X-Real datasets, it includes over 48,000 frames and 1.45 million question-answer (QA) pairs that simulate vehicle-to-vehicle communication. The QA system supports object grounding, identification, and driving planning tasks in multi-agent environments using synchronized LiDAR and camera data. Each frame contains five QA types generated through geometric templates and contextual rules. The dataset enables training and evaluation of vision-language models, like V2V-LLM, which outperform standard LLMs in cooperative perception and intention prediction. Strengths include its focus on explainability, situational awareness, and multi-agent interaction. Limitations include dependency on specific sensor setups, restricted geographic coverage, high computational demands, and uncertain usage permissions. NuScenes-QA dataset released in 2023 by Fudan University [185]. It is a large-scale VQA benchmark built on the nuScenes dataset to enhance reasoning in autonomous driving. It contains 460,000 QA pairs derived from 34,000 driving scenes with synchronized RGB images and LiDAR point clouds. Using 3D scene graphs and logic-based templates adapted from the CLEVR benchmark, researchers generated diverse questions ranging from basic attributes to complex spatiotemporal inference. Temporal questions require multi-frame reasoning to track changes and object motion. The dataset evaluates vision-language models and includes baseline results from Transformer-based models. Although NuScenes-QA advances VQA for driving tasks, limitations include a lack of background scene understanding, limited linguistic diversity due to templated questions, and shallow semantic depth.
@@
The MAPLM-QA dataset [27], introduced in 2024 through a collaboration between Tencent’s T Lab and several universities. It is a large-scale multimodal benchmark for vision-language reasoning in autonomous driving. It includes 61,000 QA pairs linked to 14,000 sensor-rich frames featuring panoramic RGB images, 3D LiDAR point clouds, and HD maps from diverse real-world environments like intersections, highways, and rural roads. To overcome limitations of prior VQA datasets, MAPLM-QA challenges LLMs to reason about road layouts, markings, and navigation cues. Data was collected using vehicles equipped with six cameras, LiDAR, and GPS/IMU, with QA annotations refined through a vision model and human verification. The dataset connects language prompts to spatial and environmental features and offers JSON-formatted tools for integration and evaluation. By evaluating LLMs like CLIP and LLaMA, this benchmark highlights their effectiveness in real-world grounding and exposes their limitations in spatial reasoning. Although it is limited by geographic scope and high computational requirements, MAPLM-QA is a pioneering resource for advancing LLM-based perception and planning. The DriveLM dataset [208], released in January 2025 by OpenDriveLab, is a large-scale benchmark for evaluating LLMs and VLMs in autonomous driving via Graph VQA. Unlike standard perception-control datasets, DriveLM emphasizes multi-step reasoning across perception, prediction, planning, and behavior tasks. It includes over 74,000 annotated frames across two subsets: DriveLM-nuScenes (4,871 frames, $\sim$91.4 QAs/frame) and DriveLM-CARLA (70,006 frames, $\sim$20.5 QAs/frame). DriveLM-nuScenes uses keyframe-based manual annotation with automatic QA generation, while DriveLM-CARLA relies on expert-driven simulation and sensor data (RGB + LiDAR). DriveLM’s strengths lie in explainability, zero-shot generalization, and structured decision flow. Limitations include reliance on simulation-based data and the high resource demands.
@@ 4.3.2 Camera-only Datasets
CODA-LM [31], NuInstruct [61], OmniDrive [238], Lingo-QA [167], and DARAMA [163] datasets rely solely on RGB visual input, often from front-facing or panoramic cameras, and are commonly used in visual-language reasoning, question answering, and instruction-following without explicit 3D spatial data. Dalian University of Technology introduced the CODA-LM dataset [31] in 2024 as a comprehensive benchmark for assessing Large Vision-Language Models (LVLMs) in practical autonomous driving systems. In order to evaluate multimodal understanding, it consists of 9,768 high-resolution scenes with uncommon driving events combined with textual prompts. CODA-LM supports three subtasks: general perception (object/event detection), regional perception (localized object description), and driving suggestion (context-based decision reasoning). It uses a hierarchical evaluation system with automated text-only LLM scoring, which offers better alignment with human judgment than multimodal evaluators. Additionally, it introduces CODA-VLM, a fine-tuned model that outperforms all open-sourced driving LVLMs and achieves GPT-4V-level results. Strengths include its focus on safety-critical corner cases, structured reasoning, and reduced reliance on human evaluation. Limitations include restricted coverage of all rare driving scenarios and high computational demands, which may limit accessibility.
@@
In 2024, HKUST, Sun Yat-Sen University, and Huawei Noah’s Ark Lab collaborated to release the NuInstruct dataset [61]. It is a large-scale instruction-based multimodal benchmark for autonomous driving. It contains more than 91,000 instruction-response pairs derived from multiview driving videos, addressing limitations in prior datasets by incorporating spatial-temporal data and diverse perspectives. The dataset is organized into 17 subtasks across four categories, including perception, prediction, risk, and planning, with reasoning to support real-world driving decision-making processes. QA pairs are generated using an SQL-based method and require reasoning across frames and views. Strengths include comprehensive coverage of driving tasks and temporal Bird’s-Eye-View cross-view reasoning. However, it lacks 3D object detection and traffic light tasks, requires high computational resources, and may need domain adaptation for generalization. OmniDrive dataset [238], released in April 2025 by Hong Kong Polytechnic University and Beijing Institute of Technology. It is a state-of-the-art multimodal dataset that links 3D driving tasks with LLMs using counterfactual reasoning. Built on nuScenes, it adds 34,000 annotated frames and 434,000 QAs focused on ”what if” scenarios that bridge perception, planning, and reasoning. The dataset includes factual and counterfactual QA pairs, trajectory data, and keyframes in JSON format for transformer-based pipelines. Two evaluation agents, Omni-Q (trajectory supervision) and Omni-L (vision-language alignment), show improved scene understanding and planning response. OmniDrive’s strengths lie in integrating perception and decision-making while enhancing interpretability and explainability for AVs. Limitations include synthetic scene generation that reduce realism and high computational demands.
@@
The Lingo-QA dataset [167], released in September 2024 by Wayve Technologies, is a large-scale multimodal benchmark for evaluating LLMs in autonomous driving. It includes 28,000 short video scenarios paired with 419,000 QA pairs covering tasks like hazard anticipation, traffic comprehension, and high-level route reasoning. Lingo-QA is built for fine-grained evaluation with QA annotations to evaluate visual perception and decision-making. Key strengths of the Lingo-QA dataset include robust annotation, multi-level reasoning, and standardized assessment. However, its limitations include reliance on short, front-facing video clips with no LiDAR or sensor data, limited scalability (up to 7B LLMs), and Lingo-Judge’s dependence on the dataset and lack of support for diverse response formats or preference-based evaluation. DRAMA dataset (Driving Risk Assessment Mechanism with A captioning module) [163], released in 2022 by Honda Research Institute USA, is a multimodal benchmark for developing risk-aware and explainable decision-making in autonomous vehicles. The DRAMA dataset combines video and object-level QA with free-form captions explaining driver reactions, linking risk localization to natural language reasoning. It features synchronized video from two front-facing cameras, along with CAN and IMU data. Annotations include open- and closed-ended questions about risk identification, causes, and reasoning, supported by natural language explanations. DRAMA’s key strength lies in integrating multimodal risk perception with interpretability for LVLMs. However, its reliance on subjective human annotations, limited coverage of failure cases, and the risks of deploying incorrect models in safety-critical settings require careful use and additional validation for real-world applications.
@@ 4.4 Vehicle-to-Vehicle (V2V) Datasets
[130] Open Mars 2024 USA +2M RGB, LiDAR, Trajectories Yes Yes No Git
[157] OPV2V-H 2024 Carla +9K RGB, LiDAR, Depth Yes Yes No Git
[94] OPV2V+ 2023 Carla 11,464 RGB, LiDAR, Depth Yes Yes No Web
@@
[269] V2V4Real 2023 USA 60K RGB, LiDAR, HDMaps Yes Yes No Web
[12] LUCOOP 2023 Germany +54K LiDAR, IMU, GNSS Yes Yes No Web
[93] CP-UAV 2022 AirSim 131.9K RGB, LiDAR, BEW Yes Yes No Git
[271] OPV2V 2022 Carla 11,464 RGB, LiDAR, BEW Yes Yes No Web
[11] CODD 2021 Carla 8,783 LiDAR, Transform.txt No Yes No Git
[298] COMAP 2021 Carla +4.3K RGB, LiDAR, Seg-Mask Yes Yes No Git
[35] T&J Cooper 2019 USA 100 RGB, LiDAR, GPS Yes Yes Yes N/A